OpenStudy (anonymous):

For x^1/2 – x^1/4 − 2 = 0, use u =

myininaya (myininaya):

Hope this helps :) \[(x^\frac{1}{4})^2-(x^\frac{1}{4})-2=0\]

OpenStudy (anonymous):

Thank you!